here's one of them: after rowing 2 miles upstream, a canoeist passes a log floating with the current. He travels 1 hour and then turns back and arrives home hust as the log gets there. Find the rate of the current.
2:the ratio of 2 posotive numbers is 3:2. The bigger number divided my the 10 less than the smaller is 2 with the remainder of 8. What is the smaller number?
3: Hallmark is going out of business! Amanda bought 6 b-day cards and 5 grad. cards for $1.52. Later in the day she decided to buy 3 b-day cards and 7 grad cards for $1.48. What is the proce for each kind of card?

Answer by rajagopalan(174)

2. Let the bigger number = x
Let the other number = y
we get x:y = 3:2
2x = 3y
x =(3/2)y
again (x-8)/(y-10) = 2 (big divided by 10 less than small) = quo 2, rem 8)
x-8 = 2(y-10)
x-8 = 2y-20
x-2y = -20+8
subst x=(3/2)y,
(3/2)y - 2y = -20+8
-(1/2)y = -12
y = 24 that is it.